OpenStudy (anonymous):

A winery has a vat with two pipes leading to it. The inlet pipe can fill the vat in 4 hours, while the outlet pipe can empty it in 10 hours. How long will it take to fill the vat if both pipes are left open?

OpenStudy (anonymous):

water filled in 1 hour by inlet pipe is 1/4 Water emptied by outlet pipe is 1/10 Total water filling the vat in 1 hour is 1/4 - 1/10 10-4/40 = 6/40 Total time taken to fill the vat is 40/6 = 6 .66 hours = 6 hours and 2/3 *60 minutes = 6 hours 40 minutes